# Are You There God? Lightweight Narrative Annotation of Christian Fiction with LMs

### v1

@@ Abstract
In addition to its more widely studied political activities, the American Evangelical movement has a well-developed but less externally visible cultural and literary side. Christian Fiction, however, has been little studied, and what scholarly attention there is has focused on the explosively popular Left Behind series. In this work, we use computational tools to provide both a broad topical overview of Christian Fiction as a genre and a more directed exploration of how its authors depict divine acts. Working with human annotators we first developed definitions and a codebook for “acts of God.” We then adapted those instructions designed for human annotators for use by a recent, lightweight LM with the assistance of a much larger model. The laptop-scale LM is capable of matching human annotations, even when the task is subtle and challenging. Using these annotations, we show that significant and meaningful differences exist between the Left Behind books and Christian Fiction more broadly and between books by male and female authors.
@@ 1 Introduction
The American Evangelical movement has sought and achieved considerable political power over the past several decades. Accordingly, there has been extensive research on Evangelicalism’s historic evolution and efforts to challenge American secularism [6, 15, 19, 31, 43, 49, 13]. However, these political movements are overwhelmingly male-dominated and therefore studies of them center the masculine Evangelical approach to asserting Christian cultural dominance [13]. Christian Fiction, a female-dominated domain [9], has received far less academic study despite having wide readership and billions of dollars in sales [42]. Christian Fiction emerged as a genre in the 1970s, establishing a Christian alternative to mainstream secular entertainment that centers a Christian worldview and values [21]. Scholarly attention to Christian Fiction has been limited and skewed towards the commercially dominant apocalyptic Left Behind series [16, 21, 46, 35, 8, 11, 30, 2].
@@
In this paper, we provide a contemporary overview of Christian Fiction using computational methodologies. We study a corpus of 88 novels: 80 that won or were short-listed for a Christy Award, offered by the Evangelical Christian Publishers Association for excellence in Christian Fiction, and the remaining 8 novels from the Left Behind series. [^1] Four novels from the 12 book series were included in our initial 80-book sample of Christy Award honorees. We begin by characterizing the main themes of these works using an established method, statistical topic modeling. We find that the collection encompasses a wide range of themes, only some of which are distinctively Christian. The Left Behind books differ significantly from the rest of the collection along several thematic axes. Moreover, even when the Left Behind books are removed from the corpus, female authors feature some traditionally feminine themes more than their male counterparts. Next, we explore a more subtle and complicated question: identifying and characterizing the portrayal of acts of God. To do this, we adapt human annotation prompts for use with a small, lightweight LM (Gemma 3n) using a much larger model (GPT-4o). Again, we find that the Left Behind novels stand out from the rest of the corpus; they feature acts of God significantly more frequently, depict more divine acts that affect groups, and are more likely to characterize those acts as punishing. Additionally, although works by male and female authors generally feature acts of God to similar extents, we see that the acts portrayed by female authors are often more loving in nature.
@@
Overall, this work shows that the frequently spotlighted Left Behind books are not representative of the general themes of Christian Fiction or its typical portrayal of divine intervention. We also show that gendered differences exist in novels written by male and female authors, even within this subgenre. Finally, we demonstrate that large, contemporary LMs can be used to effectively adapt human annotations prompts from use with smaller LMs, even for highly subtle and complex narrative phenomena.
@@ 2 Background
2.1 Christian Fiction
Initially Christian Fiction faced significant resistance within American Christian communities, where it was perceived as deviant and threatening to believers. Two primary concerns drove this opposition: the fear that fictional narratives would be dangerously conflated with the Bible and Christian doctrine – potentially confusing readers about divine truth – and the worry that fiction would divert time away from Bible reading while arousing passion and corrupting minds, particularly those of women [36, 21]. Despite this resistance, select authors such as Grace Livingston Hill began writing Christian Fiction as early as 1908 for a relatively small audience, demonstrating the genre’s persistent appeal even within hostile religious environments.
Christian Fiction gained momentum in the 1970s as American Evangelicalism emerged as a mainstream cultural movement responding to social upheaval around race, gender, and sexuality [35]. Rather than shunning mainstream culture as earlier Christian Fundamentalists had done, Evangelicals now encouraged creating alternatives to secular consumption, leading to Christian Fiction’s growth. One series in particular would come to dominate both the market and scholarly understanding of the genre: the Left Behind series by Tim LaHaye and Jerry B. Jenkins. The 2001 success of the ninth book in the series, Desecration, dethroned John Grisham’s seven-year streak as the number one bestselling author after only three months [32].
@@ 2.1 Christian Fiction
Left Behind’s success cannot be understated. LaHaye and Jenkins’ depiction of the apocalyptic Biblical Book of Revelation playing out in contemporary society sold over 80 million copies and thrust the once niche Christian Fiction industry into the mainstream [12]. The 12 book series depicts a prophetic future for the world informed by LaHaye’s literalist interpretation of the Bible [46]. [^2] LaHaye applies a premillennial dispensationalist perspective [4]. The series explores the battle against the Antichrist for control over the planet and is set within a soon approaching future, incorporating real locations, governments, and political parties. The series generated substantial academic attention that still constitutes the bulk of research on Christian Fiction. This includes discussion of the books’ perspective on Christian faith [21, 20], race [32], gender [8], and politics [35, 11]. Yet this scholarship presents significant problems: not only is the book series increasingly outdated — the final book in the Left Behind series was released over 20 years ago – but more importantly, Left Behind was co-authored by a prominent figure in Evangelical political movements, and their apocalyptic narrative contains patriarchal themes potentially unrepresentative of the novels that define contemporary Christian Fiction [8, 13]. While some scholarship has addressed more general trends in Christian Fiction [42] or women-dominated genres [36, 3], these studies remain limited in their examination of the broader Christian Fiction community. In addition, these studies focus on a very small number of texts, limiting their generalizability.
@@ 3 Data
Our corpus includes 80 Christian Fiction novels by 71 unique authors published between 2000 and 2023. These books are sampled from a comprehensive list of Christy Award book winners and finalists published between 2000 and 2024. The Christy Award was established by the Evangelical Christian Publishers Association in 1999 to celebrate faith-based novels’ impact on contemporary culture. Books honored by the Christy Award have been acknowledged as both high quality and culturally significant within the Evangelical community. By studying these books we ensure our sample represents the most influential and respected voices in Christian Fiction.
The full corpus of Christy Award honorees comprises 647 titles from 75 publishers. Over 25 years, awards have been given in 22 distinct categories ranging from ‘First Novel’ to ‘Suspense’ and ‘North American Historical.’ Approximately three-quarters of Christy Award honorees were authored only by women. [^3] Some books were co-authored by male and female authors and are excluded from this count. A breakdown of the proportion of honorees written women for each award category can be found in Appendix A.3. In our dataset, we include all Book of the Year awardees from 2014 to 2023. The remaining 70 books we sample from the 17 other award categories of interest, [^4] The three excluded categories are the Amplify Award (only given in 2022 and 2023), Lits (only given in 2007 and 2008), and Short Form (which includes novellas and short stories, but not novels). prioritizing ebook accessibility. Our corpus includes novels from 12 unique publishers, five of which account for over 61% of Christy Award honorees. [^5] These publishers are: Bethany House Publishing, Revell / Baker Publishing Group, Thomas Nelson, Tyndale House Publishers, and WaterBrook Multnomah Publishing Group As in the collection of all honoreees, three-quarters of the books in our subset were authored by women. The number of books we sampled from each award category can be found in Appendix A.2. 15% of books in our corpus won multiple Christy awards, compared to 4.5% in the entire dataset; this results from including all Book of the Year awardees, all of which won in another award category in the same year. An overview comparison of our subsample to the entire set of Christy Award honorees can be found in Table 1.
@@ 4 Methods
Our goal is to observe and quantify the thematic contents of our collection of Christian Fiction. We approach this question from both a broad, unsupervised perspective and from a more specific, targeted perspective. These goals require different methodologies.
@@ 4.2 Identifying Acts of God
For a more focused analysis, we sought to identify and characterize “acts of God” in our corpus. Though the books are fictitious, guidelines for publishing Christian Fiction [9] and authors’ unwillingness to inaccurately portray God lead to stories that represent the authors’ religious beliefs. This contributes to the blurring of the line between sacred and fictional aspects of Christian Fiction [21, 46, 16, 38]. Identifying and classifying these acts of God required both significant interdisciplinary scholarship and sophisticated contemporary language models.
@@ 4.2.1 Codebook Creation
For annotation, we split the books in our corpus into $\leq$ 500 word passages. This differs from the topic modeling data preparation because longer passages were required to provide sufficient context for the annotation process. We were left with 20,440 unique passages. On average, there are 232.27 passages per volume, with a minimum of 29 passages [^6] Candle in the Darkness by Lynn Austin and a maximum of 473 passages. [^7] The Indwelling by Jerry B. Jenkins and Tim LaHaye The passages are between 2 and 500 words long, with an average length of 430.80 words.
@@
The codebook and annotations were created by a team of ten: the lead authors with post-secondary degrees in literature, sociology, and computer science and eight undergraduates with a variety of cultural and academic backgrounds. Initially, we expected the coding task to be straightforward; the undergraduates were instructed to code a passage as “Yes” if God explicitly acted and “No” otherwise. However, the Krippendorff’s $\alpha$ scores [28] for the first two annotations rounds were very low (0.36 and 0.48). Perhaps unsurprisingly, we found that God’s presence in Christian Fiction narratives is multidimensional and often challenging to concretely identify. We also found that students’ familiarity with and relationship to Christian theology dramatically changed their interpretations of what was considered an “act” of God. To develop a unified and concrete understanding of Evangelical depictions of God, the team spent two semesters meeting weekly. Students were assigned relevant literature on American Christianity, Evangelicalism, and Christian Fiction to establish a common language and knowledge of the topic. Each week, students were tasked with coding a set of passages and asked to present on passages they found challenging.
@@
Student feedback and extensive discussions among the entire team were used to make several key changes to the codebook. First, we decided that it was important to code from an Evangelical perspective. This meant that a passage was labeled “Yes” whenever an action was attributed to God, however mundane. Therefore, we coded actions ranging from thanking God for providing a delicious meal to God sending apocalyptic plagues. Second, the adoption of the Evangelical perspective meant that the Bible was considered the literal word of God and Bible quotes or stories describing God’s actions were labeled “Yes.” Finally, we introduced a “Maybe” label to cover edge cases. These changes resulted in an increase in inter-rater reliability to $\alpha\geq 0.65$. Given the difficulty of task and the large number of annotators, we considered this sufficient.
@@
We chose the $\leq$500 word segmentation method because it was unambiguous and could be implemented quickly and accurately, but there are both technical and literary complications. In some cases, acts of God would fall across the arbitrary segmentation boundaries, making them difficult to accurately identify. There were also seeming acts of God that would later be revealed as intentional misdirection by the author. Finally, narrative-specific references to God’s actions may have been missed because the coders were not familiar with the entire texts. We do not believe, however, that these issues were common enough to threaten the overall validity of our results.
@@ 4.2.2 Annotation with LLMs
The model also repeatedly struggled to distinguish between acts of God and other supernatural or magical events, like those commonly depicted in science fiction or fantasy novels. Our attempts to clarify this distinction in the original prompt led to decreased model performance for other passages. Therefore, we introduced a second prompt to the classification pipeline that asked the model to clarify between acts of God and supernatural events. All passages labeled “Yes” by the original classifier were fed to the second prompt; passages only received the final “Yes” label if both prompts labeled them “Yes.” The second prompt went through the same iterative editing process by GPT-4o and the research team. The final versions of each prompt are available in Figure 1.
@@
On the entire evaluation dataset, the classification pipeline achieved an overall F1 score of 0.87 (Table 2). We consider this an impressive result for such a subtle and complex task. The recall for both labels is high, but we find that the model continued to over-label passages as “Yes.” Specifically, it labeled passages in which a character prayed or otherwise spoke to God as “Yes,” although the passage did not include any implied or explicit response by God and despite instructions against this in the prompts. Overall, we find that iterative editing by the lead authors and a large, powerful LM converted human annotator instructions into a successful prompt for a smaller, accessible LM.

### v2

@@ Abstract
In addition to its more widely studied cultural movements, American Evangelicalism has a well-developed but less externally visible literary side. Christian Fiction, however, has been little studied, and what scholarly attention there is has focused on the explosively popular Left Behind series. In this work, we use computational tools to provide both a broad topical overview of Christian Fiction as a genre and a more directed exploration of how its authors depict divine acts. Working with human annotators, we first developed a codebook for identifying “acts of God.” We then adapted the codebook for use by a recent, lightweight LM with the assistance of a much larger model. The laptop-scale LM is largely capable of matching human annotations, even when the task is subtle and challenging. Using these annotations, we show that significant and meaningful differences exist between divine acts depicted by the Left Behind books and Christian Fiction more broadly.
@@ 1 Introduction
The American Evangelical movement has achieved considerable cultural power over the past several decades. Accordingly, there has been extensive research on Evangelicalism’s historic evolution and efforts to challenge American secularism [6, 17, 21, 33, 45, 51, 15]. However, Evangelical ministry is overwhelmingly male-dominated and, therefore, studies of it center the masculine Evangelical approach to asserting Christian cultural dominance [15]. Christian Fiction, a female-dominated domain [10], has received far less academic study despite having wide readership and billions of dollars in sales [44]. It formally established itself as a genre in the 1970s, creating a Christian alternative to mainstream secular entertainment [23]. Christian Fiction authors occupy a unique position within American Christianity; when they depict God acting in their stories – speaking to characters, answering prayers, intervening in crises – they make consequential spiritual and creative choices about divine character and behavior. Existing scholarship on Christian Fiction has been limited and skewed towards the commercially dominant Left Behind series [18, 23, 48, 37, 9, 12, 32, 2]. Written by two men in the turn of the 20th century, this apocalyptic series does not represent what Christian Fiction is today: a women-dominated industry of primarily contemporary and historical romance, producing hundreds of titles annually that center women’s spiritual lives and everyday faith [38].
@@
In this paper, we seek to provide a contemporary overview of Christian Fiction using computational methodologies. We study a corpus of 88 novels: 80 that won or were short-listed for a Christy Award, offered by the Evangelical Christian Publishers Association for excellence in Christian Fiction, and the remaining 8 novels from the Left Behind series. [^1] Four novels from the 12 book series were included in our initial 80-book sample of Christy Award honorees. We begin by characterizing the main themes of these works using an established method, statistical topic modeling. We find that the collection encompasses a wide range of themes, only some of which are distinctively Christian. The Left Behind books differ significantly from the rest of the collection along several thematic axes. Moreover, even when the Left Behind books are removed from the corpus, themes depicted by male and female authors differ in traditionally gendered ways. Next, we identify and characterize how authors portray acts of God in their stories. To do this, we adapt human annotation prompts for use with a small, lightweight LM (Gemma 3n) using a much larger model (GPT-4o). Again, we find that the Left Behind novels stand out from the rest of the corpus; they feature acts of God significantly more frequently, depict more divine acts that affect groups, and are more likely to characterize those acts as punishing.
@@
Overall, we show that the frequently spotlighted Left Behind series is not representative of the general themes of Christian Fiction or its typical portrayal of divine intervention; instead, the authors who dominate the contemporary Christian Fiction landscape portray a loving God who overwhelmingly intercedes on the behalf of individuals. We also find some differences between contemporary male and female authors; namely, women authors focus more on joyful depcitions of affection and less on violence than their male counterparts. Finally, this paper demonstrates that large, contemporary LMs can be used to effectively adapt human annotation prompts for use with smaller LMs, even for highly subtle and complex narrative phenomena.
@@ 2 Background
2.1 Evangelicalism + Christian Fiction
Evangelicalism as an American religio-cultural movement defies simple categorization, with scholars offering a variety of overlapping yet distinct definitions [44, 17, 8, 20]. Despite this debate, Evangelicals are generally distinguished from the broader Christian community by several core beliefs: upholding the Bible as the ultimate spiritual authority that should be read and interpreted literally, experiencing a born-again conversion or moment of salvation, celebrating the centrality of Christ’s atonement, and actively working to spread their faith [15, 38, 52]. Christian Fiction specifically gained momentum in the 1970s as American Evangelicalism emerged as a mainstream cultural movement responding to social upheaval around race, gender, and sexuality [37]. Rather than shunning mainstream culture as earlier Christian Fundamentalists had done, Evangelicals increasingly encouraged the creation of alternatives from secular consumption, contributing to the growth of the genre.
Twentieth-century Evangelicalism intensified patriarchy and women’s exclusion from formal ministry even as broader American culture moved toward equality [15, 3]. However, Christian Fiction represents one of the few spaces within the Evangelical movement where women can exercise ministerial authority. Authors within this genre describe their writing as divine calling, positioning themselves as spiritual vessels using narrative to facilitate readers’ encounters with their faith. Following pioneers like Janette Oke who established the legitimacy of Christian romance, contemporary authors understand their creative process as fundamentally spiritual, involving prayer and calling for God’s guidance when developing plot and character. Readers engage with these texts as devotional practice, using fiction to maintain faith and understand God’s character [18, 40, 38, 44]. This creates “a ministry by, for, and about women” [38, p. 108] operating outside institutional structures that formally restrict women’s spiritual leadership.
Examining how authors portray God’s actions allows us to directly study their perceptions of God. Authors make consequential choices about divine behavior within their narratives: how God speaks, what motivates divine intervention, how God responds to suffering, and more [9, 14]. These choices reflect authors’ theological convictions; authors would not write God acting in ways they consider blasphemous or contrary to scripture. When an author depicts God comforting an individual, calling a nation to repentance, or remaining silent during tragedy, they therefore communicate their understanding of God to readers. These narrative patterns thus reveal how Evangelical women conceptualize and teach about God’s relationship to humanity, offering insight into women’s religious expression.
@@ 2.1 Evangelicalism + Christian Fiction
Within Christian Fiction, the Left Behind series by Tim LaHaye and Jerry B. Jenkins in particular has come to dominate both the market and the scholarly understanding of the genre. Left Behind’s success cannot be understated. It sold over 80 million copies and thrust the once niche Christian Fiction industry into the mainstream [13]. The 12 book series depicts a prophetic future for the world informed by LaHaye’s literalist interpretation of the Book of Revelation [48]. The series explores how the End Time descends upon the world as the Antichrist battles for control over the planet and is set within a soon approaching future, incorporating real locations, governments, and political parties. The series generated substantial academic attention that still constitutes the bulk of research on Christian Fiction. This includes discussion of the books’ perspective on Christian faith [23, 22], race [34], gender [9], and politics [37, 12]. Yet the singular focus of this scholarship presents a significant problem; Left Behind is not representative of Christian Fiction. The series is increasingly outdated — the final book in the Left Behind series was released over 20 years ago. More importantly, Left Behind was co-authored by a prominent figure in Evangelical political movements and its apocalyptic narrative contains patriarchal themes potentially unrepresentative of the novels that define contemporary Christian Fiction [9, 15]. While some scholarship has addressed more general trends in Christian Fiction [44] or women-dominated genres [38, 4], these studies remain limited in their examination of the broader Christian Fiction community. These studies also focus on a very small number of texts, limiting their generalizability.
@@ 3 Data
Our corpus includes 80 Christian Fiction novels by 71 authors published between 2000 and 2023. These books are sampled from a comprehensive list of Christy Award book winners and finalists published between 2000 and 2024. The Christy Award was established by the Evangelical Christian Publishers Association in 1999 to celebrate faith-based novels’ impact on contemporary culture. Books honored by the Christy Award have been acknowledged as both high quality and culturally significant within the Evangelical community. By studying these books we ensure our sample represents the most influential and respected voices in Christian Fiction.
The full corpus of Christy Award honorees comprises 647 titles from 75 publishers. Over 25 years, awards have been given in 22 distinct categories ranging from ‘First Novel’ to ‘Suspense’ and ‘North American Historical.’ Approximately three-quarters of Christy Award honorees were authored only by women. [^2] Some books were co-authored by male and female authors and are excluded from this count. A breakdown of the proportion of honorees written women for each award category can be found in Appendix A.3. In our dataset, we include all Book of the Year awardees from 2014 to 2023. The remaining 70 books we sample from the 17 other award categories of interest, [^3] The three excluded categories are the Amplify Award (only given in 2022 and 2023), Lits (only given in 2007 and 2008), and Short Form (which includes novellas and short stories, but not novels). prioritizing ebook accessibility. Our corpus includes novels from 12 publishers, five of which account for over 61% of Christy Award honorees. [^4] These publishers are: Bethany House Publishing, Revell / Baker Publishing Group, Thomas Nelson, Tyndale House Publishers, and WaterBrook Multnomah Publishing Group As in the collection of all honoreees, three-quarters of the books in our subset were authored by women. The number of books we sampled from each award category can be found in Appendix A.2. 15% of books in our corpus won multiple Christy awards, compared to 4.5% in the entire dataset, because all Book of the Year awardees by definition won another award in the same year. An overview comparison of our subsample to the entire set of Christy Award honorees can be found in Table 1.
@@ 4 Methods
Our goal is to observe and quantify the thematic contents of our collection of Christian Fiction, particularly the characterization of the divine. We approach this question from both a broad, unsupervised perspective and from a more specific, targeted perspective. These goals require different methodologies.
@@ 4.2 Identifying Acts of God
For a more focused analysis, we sought to identify and characterize acts of God in our corpus. Though the books are fictitious, guidelines for publishing Christian Fiction [10] and authors’ unwillingness to inaccurately portray God lead to stories that represent the authors’ religious beliefs. This contributes to the blurring of the line between sacred and fictional aspects of Christian Fiction [23, 48, 18, 40]. Identifying and classifying these acts of God required both significant interdisciplinary scholarship and sophisticated contemporary language models.
@@ 4.2.1 Codebook Creation
For annotation, we split the books in our corpus into $\leq$500 word passages. 500 words constitute about two average novel pages; using passages of this length allows us to provide annotators with contextually rich passages while staying within the high-performing context lengths of smaller LMs. We preserve chapter and section breaks in the novels to further increase the coherency of passages. This results in 20,440 unique passages. On average, there are 232.27 passages per volume, with a minimum of 29 passages [^5] Candle in the Darkness by Lynn Austin and a maximum of 473 passages. [^6] The Indwelling by Jerry B. Jenkins and Tim LaHaye The passages are between 2 and 500 words long, with an average length of 430.80 words.
@@
The codebook and annotations were created by a team of ten: the lead authors with degrees in literature, sociology, and computer science and eight undergraduates with a variety of religious and academic backgrounds. Initially, we expected the coding task to be straightforward; the undergraduates were instructed to code a passage as “Yes” if God explicitly acted and “No” otherwise. However, the Krippendorff’s $\alpha$ scores [30] for the first two annotations rounds were very low (0.36 and 0.48). Perhaps unsurprisingly, we found that God’s presence in Christian Fiction narratives is multidimensional and often challenging to concretely identify. We also found that students’ familiarity with and relationship to Christian theology dramatically changed their interpretations of what was considered an “act” of God. To develop a unified and concrete understanding of Evangelical depictions of God, the team spent two semesters meeting weekly. Students were assigned relevant literature on American Christianity, Evangelicalism, and Christian Fiction to establish a common language and knowledge of the topic. Each week, students were tasked with coding a set of passages and asked to present on passages they found challenging.
@@
Student feedback and extensive discussions among the entire team resulted in several key changes to the codebook. First, we decided that it was important to code from an Evangelical perspective. This meant that a passage was labeled “Yes” whenever an action was attributed to God, however mundane. Therefore, we included actions ranging from thanking God for providing a delicious meal to God sending apocalyptic plagues. Second, the adoption of the Evangelical perspective meant that the Bible was considered the literal word of God and Bible quotes or stories describing God’s actions were labeled “Yes.” Finally, we introduced a “Maybe” label to cover edge cases. These changes resulted in an increase in inter-rater reliability to $\alpha\geq 0.65$. Given the difficulty of task and the large number of annotators, we considered this sufficient.
@@
We chose the $\leq$500 word segmentation method because it was unambiguous and could be implemented quickly and accurately, but there are both technical and literary complications. In some cases, acts of God would fall across the segmentation boundaries, making them difficult to accurately identify. There were also potential acts of God that would later be revealed as intentional misdirection by the author. Since annotators only used passage-level context, these acts would still be labeled “Yes.” Finally, narrative-specific references to God’s actions may have been missed because the coders were not familiar with the entire texts. We do not believe, however, that these issues were common enough to threaten the validity of our results.
@@ 4.2.2 Annotation with LLMs
The model also repeatedly struggled to distinguish between acts of God and other supernatural or magical events, like those depicted in science fiction or fantasy novels. Our attempts to clarify this distinction in the original prompt led to decreased model performance for other passages. Therefore, we introduced a second prompt to the classification pipeline that asked the model to clarify between acts of God and supernatural events. All passages labeled “Yes” by the original classifier were fed to the second prompt; passages only received the final “Yes” label if both prompts labeled them “Yes.” The second prompt went through the same iterative editing process by GPT-4o and the research team. The final versions of each prompt are available in Figure 1.
@@
The classification pipeline achieved an overall F1 score of 0.87 on the entire dataset (Table 2). We consider this an impressive result for such a subtle and complex task. However, although the recall for both labels is high, we find that the model continued to over-label passages as “Yes.” Specifically, it often labeled passages in which a character prayed or otherwise spoke to God as “Yes,” although the passage did not include any implied or explicit response by God. For an annotation task as nuanced as this, a small LM the size of Gemma 3n is still limited in its capacity. Despite this, we find that iterative editing by the lead authors and a large, powerful LM converted human annotator instructions into a successful prompt for a smaller, accessible LM that allowed for nuanced analysis on a previously infeasible scope.